OpenStudy (anonymous):

Solve 5x – 6y = –38 3x + 4y = 0

OpenStudy (zepp):

You can use the substitution method or the reduction method to solve, I prefer the substitution method. Isolate x or y 5x – 6y = –38 6y = 5x - -38 6y = 5x+38 y = (5x+38)/6

OpenStudy (zepp):

Then you can substitue y into the first equation, leaving you with only one variable. y = (5x+38)/6 5x – 6y = –38 5x - 6((5x+38)/6)) = -38